OpenStudy (anonymous):

A spaceship passes you at a speed of 0.850c. You measure its length to be 48.2m. How long would it be when at rest?

OpenStudy (theeric):

Hello! This covers the "length contraction" phenomena. You start out knowing the relative speed between you and it. That is \(.850c\) and we'll call that \(v\). You measure its length to be \(48.2\ [m]\) when it's moving that fast. But, at relative velocity near the speed of light, it has been contracted. We'll call that contracted length \(L\). So the real length must actually be longer. The formula you must use is \[L_0=L\ \gamma\] The \(\gamma\) is the Lorentz factor, which is important in many equations about special relativity. It's much easier to look at that than what it's equal to. \(\gamma=\dfrac{1}{\sqrt{1-\dfrac{v^2}{c^2}}}\). In that, \(c\) is the speed of light, and \(v\) is the relative velocity we mentioned before. Now you have \(L_0=L\ \dfrac{1}{\sqrt{1-\dfrac{v^2}{c^2}}}\). You just have to plug the values in, now!